Assume the initial ssthresh value is 64 MSS, the advertised receiver window size in ACKs is 135 MSS unless otherwise specified, and each segment is sized one MSS. The cwnd is 5 MSS at present. Assume there is no segment loss unless otherwise specified. Provide the answers to the following questions and separate them with comma. Each question should be considered independent of others.

  1. What is the TCP window size after 34 new ACKs have been received in a row?
  2. How many new ACKs must be received continuously before the TCP window size reaches 70 MSS?
  3. After the TCP window size reaches 88 MSS, 88 new consecutive ACKs are received. The advertised window size in the last received ACK is 78 MSS. What is the TCP window size following the receipt of these new ACKs?
  4. After the TCP window size reaches 88 MSS, four duplicate ACKs are received in a row. In TCP Tahoe, what is the TCP windows size following the receipt of these duplicate ACKs?
  5. After the TCP window size reaches 88 MSS, four duplicate ACKs are received in a row. In TCP Reno, what is the TCP windows size following the receipt of these duplicate ACKs?
  6. After the TCP window size reaches 88 MSS, there is a timeout. How many new ACKs must be received continuously before the TCP window size reaches 88 MSS again?

this was the expected answer, please show work

39,458,78,1,48,2925
Answer #1

1)What is the TCP window size after 34 new ACKs have been received in a row?

ans:

TCP (Transmission Control Protocol) is a connection oriented protocol which means that we keep track of how much data has been transmitted. The sender will transmit some data and the receiver has to acknowledge it. When we don’t receive the acknowledgment in time then the sender will re-transmit the data.

TCP uses “windowing” which means that a sender will send one or more data segments and the receiver will acknowledge one or all segments. When we start a TCP connection, the hosts will use a receive buffer where we temporarily store data before the application can process it.

H2 HI Segmenti ACK 2 2 10 REDMI NOTES PRO C AL QUAD CAMERA

When the receiver sends an acknowledgment, it will tell the sender how much data it can transmit before the receiver will send an acknowledgment. We call this the window size. Basically, the window size indicates the size of the receive buffer.

Typically the TCP connection will start with a small window size and every time when there is a successful acknowledgement, the window size will increase.

2)How many new ACKs must be received continuously before the TCP window size reaches 70 MSS?

Allow W to specify a larger window size measured in categories. After that, W * MSS / RTT =
10Mbps, as packets will be reduced if the maximum transmission rate exceeds the link capacity.
So, we have W * 1500 * 8 / 0.1 = 10 * 10 ^ 6, and the W is about 84 (roof 83.3).
Since the size of the density window varies from W / 2 to W, then the average window size is
0.75W = 63 parts. The average throughput is 63 * 1500 * 8 / 0.1 = 7.56Mbps.
84/2 * 0.1 = 4.2 seconds, as the number of RTTs (required by TCP connection)
increasing window size from W / 2 to W) is provided by W / 2. Remember window size
increases one by one on each RTT.